OpenStudy (anonymous):

Please. For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 −9 + 2y −9 − 2y 9 − 2y

OpenStudy (anonymous):

−9 + 2y

OpenStudy (anonymous):

Let's see substitution method Isolate x in the first equation −x + 2y = 9 Add x on both sides 2y = 9 + x Subtract 9 on both sides -9 + 2y =x So A. −9 + 2y

OpenStudy (anonymous):

Ok I see where I messed up. Thank you.